OpenStudy (anonymous):

find the limit as x→0 of cos(mx) − cos(nx) / x^2. use L'Hospital's rule if possible

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{\cos(mx)-\cos(nx)}{x^2}\]

OpenStudy (freckles):

so can we use l'hospital?

OpenStudy (freckles):

we definitely know the bottom is 0 when plugin in 0

OpenStudy (freckles):

so only question to ask do we have 0/0

OpenStudy (freckles):

if so we can use l'hospital

OpenStudy (anonymous):

I thought yes, because it's 1-1 for the numerator, which is 0.

OpenStudy (freckles):

yep yep

OpenStudy (freckles):

now we differentiate top and differentiate bot

OpenStudy (anonymous):

I know the bottom is 2x, but I'm confused about the top.

OpenStudy (anonymous):

I found a solution online that pulls the m and the n out in front of their respective cosines, but I don't get why at all.

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{\frac{d}{dx}(\cos(mx)-\cos(nx))}{\frac{d}{dx}(x^2)} \\ =\lim_{x \rightarrow 0}\frac{\frac{d}{dx} \cos(mx)-\frac{d}{dx}\cos(nx)}{2x}\]

OpenStudy (freckles):

well first of all you know the derivative of the outsides that is what is the derivative of cos

OpenStudy (anonymous):

Should I be treating cos(mx) and cos(nx) as a product rule?

OpenStudy (freckles):

well it isn't a product it is a difference

OpenStudy (freckles):

\[\frac{d}{dx}(f-g)=\frac{d}{dx}f-\frac{d}{dx}g\]

OpenStudy (freckles):

so we use difference rule

OpenStudy (freckles):

but anyways to differentiate cos(mx) we need to use chain rule

OpenStudy (freckles):

what is the derivative of cos

OpenStudy (anonymous):

-sin is the derivative of cos

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{\frac{d}{dx}(\cos(mx)-\cos(nx))}{\frac{d}{dx}(x^2)} \\ =\lim_{x \rightarrow 0}\frac{\frac{d}{dx} \cos(mx)-\frac{d}{dx}\cos(nx)}{2x} \\ =\lim_{x \rightarrow 0}\frac{(mx)'(-\sin(mx))-(nx)'(-\sin(nx))}{2x}\] right and we also need to multiply by the derivative of the inside

OpenStudy (freckles):

\[(f(g(x))'=f'(x)|_{x=g(x)} \cdot g'(x) \\ (\cos(mx))'=(\cos(x))'|_{x=mx} \cdot (mx)'\]

OpenStudy (freckles):

\[(f(g(x))'=f'(x)|_{x=g(x)} \cdot g'(x) \\ (\cos(mx))'=(\cos(x))'|_{x=mx} \cdot (mx)' \\ (\cos(mx))'=-\sin(x)|_{x=mx} (m) \\ (\cos(mx))'=-\sin(mx) (m) \\ (\cos(mx))'=-m \sin(mx)\]

OpenStudy (anonymous):

Okay, I understand that completely.

OpenStudy (freckles):

really ? if so that is great

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{\frac{d}{dx}(\cos(mx)-\cos(nx))}{\frac{d}{dx}(x^2)} \\ =\lim_{x \rightarrow 0}\frac{\frac{d}{dx} \cos(mx)-\frac{d}{dx}\cos(nx)}{2x} \\ =\lim_{x \rightarrow 0}\frac{(mx)'(-\sin(mx))-(nx)'(-\sin(nx))}{2x} \\ =\lim_{x \rightarrow 0}\frac{-m \sin(mx)+n \sin(nx)}{2x}\] so what happens if we enter in 0 for x?

OpenStudy (anonymous):

I believe you get -m * 0 + n * 0 = 0 for the numerator and 0 for the denominator, so you get 0/0 again. Which means you have to do L'Hospital's rule again

OpenStudy (freckles):

don't get frustrated this will be the last round of l'hospital because the bottom will be 2

OpenStudy (freckles):

so we only have to differentiate that nasty top one more time

OpenStudy (freckles):

our new top anyways

OpenStudy (freckles):

we need to apply the chain rule again

OpenStudy (anonymous):

hmmm, okay

OpenStudy (freckles):

do you want to try

OpenStudy (freckles):

just differentiate one term

OpenStudy (anonymous):

I will try.

OpenStudy (freckles):

\[\frac{d}{dx}(-msin(mx))=\]

OpenStudy (freckles):

that -m in front is a constant multiply you can drag it outside the derivative operator

OpenStudy (freckles):

\[\frac{d}{dx}\sin(mx) \text{ put main focus here }\] we will multiply -m to the derivative later

OpenStudy (anonymous):

So you'd get\[-m \cos(mx)*m\] ?

OpenStudy (anonymous):

so it'd be \[-m ^{2}\cos(mx)\]

OpenStudy (freckles):

\[\frac{d}{dx}(-m \sin(mx)) \\ -m \frac{d}{dx}\sin(mx) \\ -m \cos(mx) \cdot m \\ -m^2 \cos(mx) \] AWESOME! :)

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{\frac{d}{dx}(\cos(mx)-\cos(nx))}{\frac{d}{dx}(x^2)} \\ =\lim_{x \rightarrow 0}\frac{\frac{d}{dx} \cos(mx)-\frac{d}{dx}\cos(nx)}{2x} \\ =\lim_{x \rightarrow 0}\frac{(mx)'(-\sin(mx))-(nx)'(-\sin(nx))}{2x} \\ =\lim_{x \rightarrow 0}\frac{-m \sin(mx)+n \sin(nx)}{2x} \\ \lim_{x \rightarrow 0} \frac{-m^2 \cos(mx)+n^2 \cos(nx)}{2}\]

OpenStudy (freckles):

you are ready to just plug in 0

OpenStudy (anonymous):

YAY!!! So, cos(0) = 1, which means the final answer is \[\frac{ n ^{2}-m ^{2} }{ 2}\]

OpenStudy (freckles):

yah

OpenStudy (freckles):

Do you feel like anything here needs to be discussed more?

OpenStudy (anonymous):

Thank you so much for your help! I'm set for now!

OpenStudy (freckles):

Ok cool have a nice night

OpenStudy (anonymous):

You too :)